OpenStudy (anonymous):

HELP PLEASE help will medal simplify the polynomial a2+5a-3a2

OpenStudy (henriquerb):

a2 is a squared?

OpenStudy (anonymous):

yes

OpenStudy (henriquerb):

So a^2 +5a - 3a^2 = 5a - 2a^2 = a(5-2a)

OpenStudy (anonymous):

can you help me with two more? @henriquerb

OpenStudy (henriquerb):

I can try =)

OpenStudy (anonymous):

-4x+6x^2+5x

OpenStudy (anonymous):

@henriquerb

OpenStudy (henriquerb):

ok, so you'll have to add 5x with -4x. that gives x -4x+6x²+5x= 6x² +x(5-4)= 6x² +x Then, take x "out" x(6x+1)

OpenStudy (anonymous):

6g+8k-9g+2k

OpenStudy (anonymous):

@henriquerb

OpenStudy (henriquerb):

Same thing, add the g's and the k's 6g+(-9g)+8k+2k = -3g +10k

OpenStudy (henriquerb):

Are you understanding what I'm doing?

OpenStudy (anonymous):

yes and thanks for the help

OpenStudy (henriquerb):

Ok, very good. You're welcome. Anytime you need help, I'll be glad to try.